three capacitors c1 c2 and c3 are connected in series and another capacitor c4 connected in parallel in c2 and c4 are equal to 5coulomb whiles c1and c3 have a capacitance of 12coulombs. calcute the magnitude of the charge c4


1
Expert's answer
2021-08-26T08:10:54-0400
"q_4=q-q_2\\q_4=12-5=7\ C"